OpenStudy (anonymous):

eigenvalue/vector help needed! X'(t)=(1 0 0 0 1 2 0 0 1 2 -1 1 1 1 0 4) X(t) A) find the eigenvalues for the above system of equations B) find the eigenvectors for the above system of equations C) Show that answers A and B are correct.

OpenStudy (turingtest):

I would take the determinant using the fact that the top row is almost all zero's. I assume you know how to apply the eigenvalue part to the procedure.

OpenStudy (anonymous):

That's the thing... My book didn't explain it very well at all, so I have very little knowledge about how to do this... sorry. I'm hoping someone here could teach me better then my book

OpenStudy (turingtest):

Do you know the method of cofactors?

OpenStudy (anonymous):

no sorry

OpenStudy (turingtest):

well taking this determinant will be a real pain without it, so I suggest you read up on that a bit http://www.mathwords.com/e/expansion_by_cofactors.htm http://www.purplemath.com/modules/minors.htm

OpenStudy (anonymous):

oh I guess I do know that just not for a 4x4 matrix. I'll look at it some more though

OpenStudy (turingtest):

And eigenvalue \(\lambda\) is a scalar such that\[A=\lambda I\implies \det(A-\lambda I)=0\]so rewrite the matrix as \(A-\lambda I\), use the method of cofactors expanding along the top row to take the determinant, then solve for lambda to get the eigenvalues.

OpenStudy (turingtest):

oh I dropped the vector from my definition, sorry

OpenStudy (turingtest):

\[A\vec x=\lambda I\vec x\implies \det(A-\lambda I)=0\]

OpenStudy (anonymous):

okay so I start by getting the determinant of my first vector, right?

OpenStudy (turingtest):

well you start by taking the determinant of \(A-\lambda I\)

OpenStudy (anonymous):

with 'I' being? sorry I'm a bit confused

OpenStudy (turingtest):

that would be the identity matrix; a matrix with nothing but 1's on its diagonal

OpenStudy (turingtest):

\[A-\lambda I=\left[\begin{matrix}1&0&0&0\\1&2&0&0\\2&2&-1&1\\1&1&0&4\end{matrix}\right]-\lambda\left[\begin{matrix}1&0&0&0\\0&1&0&0\\0&0&1&0\\0&0&0&1\end{matrix}\right]\]

OpenStudy (turingtest):

\[A-\lambda I=\left[\begin{matrix}1-\lambda&0&0&0\\1&2-\lambda&0&0\\2&2&-1-\lambda&1\\1&1&0&4-\lambda\end{matrix}\right]\]since \(A\vec x=\lambda\vec x\) the matrices must be linearly dependent, i.e. the determinant of the above is zero

OpenStudy (turingtest):

\[\det(A-\lambda I)=\left|\begin{matrix}1-\lambda&0&0&0\\1&2-\lambda&0&0\\2&2&-1-\lambda&1\\1&1&0&4-\lambda\end{matrix}\right|=0\]expanding along the first row, this becomes...

OpenStudy (turingtest):

\[\det(A-\lambda I)=(1-\lambda)\left|\begin{matrix}2-\lambda&0&0\\2&-1-\lambda&1\\1&0&4-\lambda\end{matrix}\right|=0\]since the rest of the expansion by cofactors have coefficient zero along row 1. Take this determinant and solve for the four values of lambda, i.e. the eigenvalues.

OpenStudy (anonymous):

@TuringTest so then to find the eigenvector, I ttake the initial matrix minus one of the eigenvalues times the identity matrix, right?

OpenStudy (turingtest):

well you don't subtract just *one* of the eigenvalues, they are all encompassed in the \(\lambda\). After taking the determinant you will get a 4th order equation in terms of lambda, from which you will get 4 solutions \(\lambda_1, \lambda_2. \lambda_3, \lambda_4\), i.e. your 4 eigenvalues.

OpenStudy (anonymous):

right. I've got four eigenvalues but now I need the vectors. How do I get an answer as a vector?

OpenStudy (turingtest):

use the definition of the eigenvalue and go back to the original equation. say one of your eigenvalues is 2, you then get\[A\vec b=6\vec b\] where\[\vec b=\left[\begin{matrix}x\\y\end{matrix}\right]\]you can then solve this system for x and y, which is one of your eigenvectors

OpenStudy (turingtest):

oops, I meant "say one of your eigenvalues is 6..."

OpenStudy (turingtest):

notice that we can subtract from both sides, in which case we wind up with the same thing we had before to find the polynomial for the eigenvalues\[(A-6I)\vec b=\vec 0\]so we would just solve\[\left[\begin{matrix}-5&0&0&0\\1&-4&0&0\\2&2&-7&1\\1&1&0&-2\end{matrix}\right]\vec b=\vec 0\]

OpenStudy (anonymous):

@AccessDenied do you know anything about eigenvectors? Do you think you could help me with the eigenvector of 2 for this one? I think it's supposed to be (0,-2,-1,1) but I can't seem to get that for an answer

OpenStudy (turingtest):

did you follow what I did above with the method of cofactors?

OpenStudy (anonymous):

yep. I got it figured out but I posted another related question if you could check it out quick. Thanks for the help either way